OpenStudy (hirokocastillo):

find the equation of the line with y-intercept -4 and a slope of (3/2)

OpenStudy (luigi0210):

Do you know how to start this off?

OpenStudy (hirokocastillo):

no please help i have no clue what to do

OpenStudy (skullpatrol):

$$\huge y=mx+b$$

OpenStudy (luigi0210):

I think skull wants to help you with this

OpenStudy (hirokocastillo):

okay thank you

OpenStudy (skullpatrol):

Have you seen the form y = mx + b before?

OpenStudy (hirokocastillo):

yes but i dont know what to do with it i know mx is the slope and b is the y intercept but im still confused as to what i do with it

OpenStudy (skullpatrol):

"mx" is not the slope. Only m is the slope.

OpenStudy (hirokocastillo):

oh okay but what do i do with it

OpenStudy (skullpatrol):

find the equation of the line with y-intercept: b = -4 and a slope of m = (3/2)

OpenStudy (skullpatrol):

Does that help?

OpenStudy (hirokocastillo):

yes very much thank you so much

OpenStudy (hirokocastillo):

bye

OpenStudy (skullpatrol):

Thanks for trying to learn something new